OpenStudy (anonymous):

Write a rule for the linear function whose graph has slope m and contains the given point. m = 3; (-2, -2)

OpenStudy (anonymous):

\[y-y_1=m(x-x_1)\] with \[m=3,x_1=-2,y_1=-2\]

OpenStudy (anonymous):

first step is \[y+2=3(x+2)\]